What is the pH of a .75M H2C6H6O6? Ka1=8e-5, ka2=1.6e-12. Please show all work in ICE Tables. Thank you!
Answer #1


      H2C6H6O6 --------------------> HC6H6O6- + H+

I         0.75                                        0                   0

C          -x                                        +x                 +x

E       (0.75-x)                                   x                   x

Ka1 = [HC6H6O6-][H+]

8.0*10-5 = x2 / (0.75-x)

8.0*10-5 * (0.75-x) = x2

solving for x = 0.0078 M

so [HC6H6O6-] = [H+] = 0.0078 M

pH = -log[H+]

pH = -log[0.0078]

pH = 2.10

answer => pH = 2.1